OpenStudy (peterg2001):

help will medal Which of the following equations could you solve by first adding six and then dividing by negative three? 3x - 6 = -9 6 - 3x = -9 -3x - 6 = -9 x/-3 + 6 = -9

OpenStudy (peterg2001):

@KamiBug @shinebrightlikeadimon

OpenStudy (anonymous):

C

OpenStudy (kamibug):

Choose the one with the opposite operations on the side the variable is on. :) Which option has subtraction of 6 and multiplication of -3? :)

OpenStudy (anonymous):

-3x - 6 = -9 + 6 + 6 -3x = -3 /-3 /-3 x = 1

OpenStudy (peterg2001):

a b c or d please

OpenStudy (anonymous):

i told you C